OpenStudy (anonymous):

(2,3)(4,-5) what do you plug in?? anybody??

OpenStudy (texaschic101):

what are you trying to do ?

OpenStudy (anonymous):

she wants me to use the formulas like x1-y1 things like that and shes telling me to plug it in what do she mean??

OpenStudy (texaschic101):

(2,3) -- x1 = 2, y1 = 3 (4,-5) -- x2 = 4, y2 = -5 I don't know what else to tell you because I am not sure what she wants you to do. Does she want you to find the slope ?

OpenStudy (anonymous):

yeah shes telling me to find the slope then plug it in

OpenStudy (texaschic101):

ok....lets do this the slope formula is : slope(m) = (y2 - y1) / (x2 - x1) (2,3) - x1 = 2, y1 = 3 (4,-5) - x2 = 4, y2 = -5 now she wants you to plug the numbers into the formula can you do that ?

OpenStudy (texaschic101):

Let me start you off m = (-5 - 3) / (4 - 2)

OpenStudy (anonymous):

does that equal to -8/2

OpenStudy (texaschic101):

yes....can you reduce it ?

OpenStudy (anonymous):

hmm if I reduce it it'll be 2/1 right??

OpenStudy (texaschic101):

no.....it will be -4

OpenStudy (texaschic101):

do you understand ?

OpenStudy (anonymous):

im still trying to see how you got -4

OpenStudy (texaschic101):

-8/2......." / " means divide. How many times will 2 go into 8 ? It will go in 4 times....and don't forget your negative sign...so it is -4. Does that help ?

OpenStudy (anonymous):

o i get it now. she wants me to explain how i get the numbers and i don't know what to say that's whats failing me

OpenStudy (texaschic101):

you mean the slope ? or x 1, y 1 numbers ? ....which numbers

OpenStudy (texaschic101):

hello ???

OpenStudy (anonymous):

im here

OpenStudy (texaschic101):

which numbers are you talking about.....the slope or just the x and y numbers ?

OpenStudy (anonymous):

she tells me how did i plug in the numbers 2 points i think she jus messing with my head if u ask me

OpenStudy (texaschic101):

well...if you have 2 sets of numbers, like in your previous question.... set 1 (1,2)....set 2(2,3)....you label them (1,2) -- x1 = 1, and y 1 = 2 (2,3)--x2 = 2, and y2 = 3 it is not that hard, but with a lot of numbers it can get confusing. As far as finding the slope or if you are looking for a line, then you have to know the formulas and be able to sub the numbers in.

OpenStudy (anonymous):

yeah and that can be the only problem the y1 and y2's i just don't know how to set it up unless somebody speaks on it

OpenStudy (texaschic101):

ok.....you have 2 sets of numbers (1,2) and (3,4) set 1 will contain x1 and y1 because they come from set 1 set 2 will contain x2 and y2 because they come from set 2 does that help ? Just label your sets if that makes it easier ... set 1 (x1,y1) set 2 (x2,y2)